# Yours or Mine? Overwriting Attacks against Neural Audio Watermarking

### v1

@@ Introduction
To improve usability, existing neural audio watermarking methods primarily focus on two properties: robustness and imperceptibility. Robustness ensures that the watermark remains detectable after common signal processing operations such as compression. Imperceptibility guarantees that watermark embedding does not degrade perceptual audio quality. These objectives have driven recent progress in watermarking research. AudioSeal (San Roman et al. 2024) is based on the Encodec (Défossez et al. 2022) architecture, embedding redundant watermark messages in the embedding layer to improve robustness, while introducing novel perceptual loss to preserve audio quality. Timbre (Liu et al. 2024a) embeds the watermark in the frequency domain and incorporates a distortion layer during training to simulate real-world perturbations, thereby maintaining robustness. WavMark (Chen et al. 2023) adopts invertible neural networks (Dinh, Krueger, and Bengio 2014) to embed imperceptible watermarks and use pattern bits to improve robustness against distortions.
@@
Although the robustness of audio watermarking has been extensively studied (Wen et al. 2025; O’Reilly et al. 2024; Özer et al. 2025), its security aspects remain underexplored. Robustness refers to the ability to withstand unintentional perturbations, whereas security concerns its resilience against intentional manipulation by adversaries (Hartung and Kutter 2002; Furon and Duhamel 2003). San Roman et al. (2024) was the first to highlight potential security threats in neural audio watermarking. It showed that adversaries, with access to the watermark detector, can launch two adversarial attacks: removal attacks, which make the watermark undetectable, and forgery attacks, which falsely embed a watermark into clean audio. Liu et al. (2024b) extended these attack paradigms and systematically evaluated the vulnerabilities of neural audio watermarking methods under such threats. Furthermore, Liu et al. (2024a) discussed overwriting attacks, demonstrating that an attacker with access to the watermark embedder can insert a new watermark that effectively overwrites the original one. However, their attack relies on a white-box assumption, where the adversary has full knowledge of the watermarking framework, and the evaluation of the overwriting attack lacks thorough analysis.
@@
In this work, we present the first systematic study of watermark overwriting attacks, a previously underexplored but practically powerful threat. Unlike removal or forgery attacks, overwriting attacks embed a new watermark to replace the original legitimate one and thus hijack the ownership of the target audio. We perform comprehensive evaluations of three state-of-the-art neural audio watermarking methods (San Roman et al. 2024; Liu et al. 2024a; Chen et al. 2023) under the proposed white-box, gray-box, and black-box overwriting attacks. In all threat settings, we achieve a nearly 100% attack success rate in terms of overwriting the original watermark. These findings expose widespread vulnerabilities in existing neural audio watermarking systems and underscore the need to consider security as a primary design objective, alongside robustness and imperceptibility. Our main contributions are summarized as follows.
@@
Audio watermarking is a technique that embeds information into audio signals without significantly degrading their perceptual quality. A typical audio watermarking system consists of two key components: a watermark embedder, which encodes the information into the audio, and a watermark detector, which detects and recovers the embedded message. The watermark is typically a fixed-length binary sequence and can carry various types of information depending on the applications. For example, when the watermark encodes copyright metadata, it can support copyright declarations and infringement tracking (Liu et al. 2024a); when it includes source-related tags (e.g., indicators of AI-generated content), it enables source verification (San Roman et al. 2024); and when it contains a hash of the audio, it can be used for integrity verification and tampering detection (Yao et al. 2025).
As a proactive protection mechanism, audio watermarking systems typically require the following three properties (Hartung and Kutter 2002): robustness, imperceptibility, and security. Robustness refers to the system’s ability to retain the watermark after undergoing common audio processing or transmission (e.g., resampling, compression, or reverberation). Imperceptibility requires that the watermark embedding process introduces no perceptible distortion to audio. Security emphasizes resilience against intentional attacks, ensuring the watermark’s integrity even when adversaries attempt to manipulate it.
@@
Audio watermarking has evolved from traditional signal processing techniques into deep learning-based systems, achieving significant progress. Traditional methods, such as least significant bit (LSB) (Cvejic and Seppanen 2004), echo hiding (Gruhl, Lu, and Bender 1996), spread spectrum (Cox et al. 1997), patchwork (Yeo and Kim 2003), and quantization index modulation (QIM) (Chen and Wornell 2002), typically rely on expert knowledge and fixed rules. They are hard to implement and struggle with robustness against complex distortions.
@@
  • Encoder-decoder-based approaches, such as AudioSeal (San Roman et al. 2024), XattnMark (Liu et al. 2025), and SilentCipher (Singh et al. 2024), embed watermarks in the high-dimensional latent space learned by neural networks;
  • Frequency-domain-based approaches, such as Timbre (Liu et al. 2024a) and DeAR (Liu et al. 2023), embed watermarks into the frequency spectrum of audio;
  • Invertible neural network (INN)-based approaches, such as WavMark (Chen et al. 2023) and IDEAW (Li et al. 2024), model the embedding and detection processes as reversible transformations, enabling high-fidelity and accurate watermark recovery.
@@
Most existing neural audio watermarking methods prioritize robustness and imperceptibility, while largely overlooking security. Traditional audio watermarking methods (Furon and Duhamel 2003) often leverage the secret key to control the embedding location and detection process, thus preventing unauthorized manipulation. In contrast, neural watermarking systems typically lack such explicit key-based security mechanisms, but rely on the assumption that the secret of the model weights. However, this assumption is fragile in modern research environments, where open-sourcing and reverse engineering are common practices. According to Kerckhoffs’s principle (Kerckhoffs 1883), a secure system should remain secure even if everything about the system is public except the secret key. Therefore, relying on “security through obscurity” is inadequate for neural watermarking systems.
Recent studies have begun to explore the security vulnerabilities of neural watermarking. San Roman et al. (2024) first demonstrated that adversaries with access to the watermark detector could launch two types of adversarial attacks: removal attacks, which make the watermark undetectable, and forgery attacks, which falsely embed a watermark into clean audio. Liu et al. (2024b) extended this line of work with a more systematic evaluation of neural audio watermarking methods under adversarial attacks.
However, overwriting attacks, where adversaries embed a new watermark into an already watermarked audio to override the original ownership, remain largely underexplored. This attack poses a realistic and severe threat to ownership verification. Although Liu et al. (2024a) pointed out overwriting attacks in white-box scenarios, their discussion was limited to insider threats and did not provide a comprehensive evaluation across different adversarial settings. In this work, we address this gap by systematically investigating overwriting attacks across various threat models, detailing their implementations and evaluating their effects on existing neural audio watermarking systems.
@@
  • Frequency-domain Reconstruction Loss ($\mathcal{L}_{\text{recon}_{f}}$): To further reduce perceptual differences in the frequency domain, we introduce a multi-resolution short-time Fourier transform (STFT) loss (Yamamoto, Song, and Kim 2020). For each resolution $m$, the loss includes a spectral convergence and a log-magnitude term:
@@
To evaluate the vulnerability of neural audio watermarking systems to overwriting attacks, we conduct extensive experiments on three representative methods: AudioSeal (San Roman et al. 2024), Timbre (Liu et al. 2024a), and WavMark (Chen et al. 2023). Evaluations are performed under three threat models: white-box, gray-box, and black-box.
@@
We conduct experiments on two widely used speech datasets: LibriSpeech (Panayotov et al. 2015), a corpus of approximately 1,000 hours of English read speech, and VoxCeleb1 (Nagraniy, Chungy, and Zisserman 2017), which contains over 150,000 utterances from 1,251 celebrities. All audio samples are resampled to 16kHz and converted to WAV format for consistency across models. All models are trained on a server equipped with 64 GPU cores and two NVIDIA A100 GPUs.
@@
The diagonal entries (the same method used for both original and overwriting watermarks) consistently reach a BER near 0.5, which means random guessing. This confirms that the original watermark is completely disrupted when the same watermarking method is reused for overwriting. In contrast, off-diagonal values (different methods) show very low BERs, suggesting that overwriting using a different watermarking method fails to destroy the original legitimate watermark. This is because different watermarking methods operate in distinct embedding domains and rely on method-specific decoding mechanisms. For example, AudioSeal embeds watermarks as additive perturbations in the Encodec (Défossez et al. 2022) latent space, while Timbre modifies frequency-domain representations. Therefore, their watermarks do not interfere with each other.
@@
Metric Timbre AudioSeal WavMark
ASR (Original Watermark) 99.80 100.00 100.00
ACC (Overwritten Watermark) 100.00 100.00 100.00
@@
In the gray-box setting, the adversary knows the watermarking system’s architecture but lacks access to its training details or training data. To assess this scenario, we construct surrogate models under two settings: (i) cross-training, where the surrogate model is trained on the same dataset (VoxCeleb1) but with a different training pipeline, and (ii) cross-data, where the surrogate is trained on an entirely different dataset (LibriSpeech). We train three surrogate models with varying random seeds (42, 99, and 2025), denoted as Init-1, Init-2, and Init-3, respectively. The “Init-$k\rightarrowOfficial$” notation indicates that surrogate Init-$k$ attempts to overwrite the watermark embedded by the official model.
@@
Method (LibriSpeech $\rightarrow$ VoxCeleb1) Init-1 Init-2 Init-3
@@
The query-based attack mitigates these drawbacks by utilizing limited queries to the watermark detector. Instead of exhaustively applying all candidates, the adversary iteratively “try-and-test” candidates and stops once the detector confirms successful overwriting of the original legitimate watermark. Consequently, audio degradation is limited to a single embedding operation. As depicted in Figure 5, watermarking models exhibit varying convergence time. Certain methods reach sufficient overwriting capabilities early in the training. This property enables successful overwriting using partially trained embedders, significantly reducing the required training iterations.
Table 4 provides a comparative analysis between zero-query and query-based attacks. The results indicate that query-based attacks, using fewer than 10 detector queries, achieve a reduction of over 50% in training iterations compared to the zero-query approach. By applying only a single effective watermarking method instead of stacking multiple methods, query-based attacks preserve the audio quality and maintain identical attack success rates. As the size of the candidate set grows, the advantages of the query-based approach become more pronounced.
@@
Attack Type Query Count Training Cost SNR (dB) ASR (%)
Zero-query 0 36,000 iters 24.19 100
Query-based <<10 14,000 iters 20.63 100

### v2

@@ Introduction
To improve usability, existing neural audio watermarking methods primarily focus on two properties: robustness and imperceptibility. Robustness ensures that the watermark remains detectable after common signal processing operations such as compression. Imperceptibility guarantees that watermark embedding does not degrade perceptual audio quality. These objectives have driven recent progress in watermarking research. AudioSeal (san2024proactive) is based on the Encodec (defossez2022high) architecture, embedding redundant watermark messages in the embedding layer to improve robustness, while introducing novel perceptual loss to preserve audio quality. Timbre (liu2024detecting) embeds the watermark in the frequency domain and incorporates a distortion layer during training to simulate real-world perturbations, thereby maintaining robustness. WavMark (chen2023wavmark) adopts invertible neural networks (dinh2014nice) to embed imperceptible watermarks and use pattern bits to improve robustness against distortions.
@@
Although the robustness of audio watermarking has been extensively studied (wen2025sok; shallowwatermark; ozer2025comprehensive), its security aspects remain underexplored. Robustness refers to the ability to withstand unintentional perturbations, whereas security concerns its resilience against intentional manipulation by adversaries (hartung2002multimedia; furon2003asymmetric; li2021talk). san2024proactive was the first to highlight potential security threats in neural audio watermarking. It showed that adversaries, with access to the watermark detector, can launch two adversarial attacks: removal attacks, which make the watermark undetectable, and forgery attacks, which falsely embed a watermark into clean audio. liu2024audiomarkbench extended these attack paradigms and systematically evaluated the vulnerabilities of neural audio watermarking methods under such threats. Furthermore, liu2024detecting discussed overwriting attacks, demonstrating that an attacker with access to the watermark embedder can insert a new watermark that effectively overwrites the original one. However, their attack relies on a white-box assumption, where the adversary has full knowledge of the watermarking framework, and the evaluation of the overwriting attack lacks thorough analysis.
@@
In this work, we present the first systematic study of watermark overwriting attacks, a previously underexplored but practically powerful threat. Unlike removal or forgery attacks, overwriting attacks embed a new watermark to replace the original legitimate one and thus hijack the ownership of the target audio. We perform comprehensive evaluations of three state-of-the-art neural audio watermarking methods (san2024proactive; liu2024detecting; chen2023wavmark) under the proposed white-box, gray-box, and black-box overwriting attacks. In all threat settings, we achieve a nearly 100% attack success rate in terms of overwriting the original watermark. These findings expose widespread vulnerabilities in existing neural audio watermarking systems and underscore the need to consider security as a primary design objective, alongside robustness and imperceptibility. Our main contributions are summarized as follows.
@@
Audio watermarking is a technique that embeds information into audio signals without significantly degrading their perceptual quality. A typical audio watermarking system consists of two key components: a watermark embedder, which encodes the information into the audio, and a watermark detector, which detects and recovers the embedded message. The watermark is typically a fixed-length binary sequence and can carry various types of information depending on the applications. For example, when the watermark encodes copyright metadata, it can support copyright declarations and infringement tracking (liu2024detecting); when it includes source-related tags (e.g., indicators of AI-generated content), it enables source verification (san2024proactive); and when it contains a hash of the audio, it can be used for integrity verification and tampering detection (yao2025speechverifier).
As a proactive protection mechanism, audio watermarking systems typically require the following three properties (hartung2002multimedia): robustness, imperceptibility, and security. Robustness refers to the system’s ability to retain the watermark after undergoing common audio processing or transmission (e.g., resampling, compression, or reverberation). Imperceptibility requires that the watermark embedding process introduces no perceptible distortion to audio. Security emphasizes resilience against intentional attacks, ensuring the watermark’s integrity even when adversaries attempt to manipulate it.
@@
Audio watermarking has evolved from traditional signal processing techniques into deep learning-based systems, achieving significant progress. Traditional methods, such as least significant bit (LSB) (cvejic2004increasing), echo hiding (gruhl1996echo), spread spectrum (cox1997secure), patchwork (yeo2003modified), and quantization index modulation (QIM) (chen2002quantization), typically rely on expert knowledge and fixed rules. They are hard to implement and struggle with robustness against complex distortions.
@@
  • Encoder-decoder-based approaches, such as AudioSeal (san2024proactive), XattnMark (liu2025xattnmark), and SilentCipher (singh2024silentcipher), embed watermarks in the high-dimensional latent space learned by neural networks;
  • Frequency-domain-based approaches, such as Timbre (liu2024detecting) and DeAR (liu2023dear), embed watermarks into the frequency spectrum of audio;
  • Invertible neural network (INN)-based approaches, such as WavMark (chen2023wavmark) and IDEAW (li2024ideaw), model the embedding and detection processes as reversible transformations, enabling high-fidelity and accurate watermark recovery.
@@
Most existing neural audio watermarking methods prioritize robustness and imperceptibility, while largely overlooking security. Traditional audio watermarking methods (furon2003asymmetric) often leverage the secret key to control the embedding location and detection process, thus preventing unauthorized manipulation. In contrast, neural watermarking systems typically lack such explicit key-based security mechanisms, but rely on the assumption that the secret of the model weights. However, this assumption is fragile in modern research environments, where open-sourcing and reverse engineering are common practices. According to Kerckhoffs’s principle (kerckhoffs1883cryptographie), a secure system should remain secure even if everything about the system is public except the secret key. Therefore, relying on “security through obscurity” is inadequate for neural watermarking systems.
Recent studies have begun to explore the security vulnerabilities of neural watermarking. san2024proactive first demonstrated that adversaries with access to the watermark detector could launch two types of adversarial attacks: removal attacks, which make the watermark undetectable, and forgery attacks, which falsely embed a watermark into clean audio. liu2024audiomarkbench extended this line of work with a more systematic evaluation of neural audio watermarking methods under adversarial attacks.
However, overwriting attacks, where adversaries embed a new watermark into an already watermarked audio to override the original ownership, remain largely underexplored. This attack poses a realistic and severe threat to ownership verification. Although liu2024detecting pointed out overwriting attacks in white-box scenarios, their discussion was limited to insider threats and did not provide a comprehensive evaluation across different adversarial settings. In this work, we address this gap by systematically investigating overwriting attacks across various threat models, detailing their implementations and evaluating their effects on existing neural audio watermarking systems.
@@
  • Frequency-domain Reconstruction Loss ($\mathcal{L}_{\text{recon}_{f}}$): To further reduce perceptual differences in the frequency domain, we introduce a multi-resolution short-time Fourier transform (STFT) loss (yamamoto2020parallel). For each resolution $m$, the loss includes a spectral convergence and a log-magnitude term:
@@
To evaluate the vulnerability of neural audio watermarking systems to overwriting attacks, we conduct extensive experiments on three representative methods: AudioSeal (san2024proactive), Timbre (liu2024detecting), and WavMark (chen2023wavmark). Evaluations are performed under three threat models: white-box, gray-box, and black-box.
@@
We conduct experiments on two widely used speech datasets: LibriSpeech (panayotov2015librispeech), a corpus of approximately 1,000 hours of English read speech, and VoxCeleb1 (nagraniy2017voxceleb), which contains over 150,000 utterances from 1,251 celebrities. All audio samples are resampled to 16kHz and converted to WAV format for consistency across models. All models are trained on a server equipped with 64 CPU cores and two NVIDIA A100 GPUs.
@@
The diagonal entries (the same method used for both original and overwriting watermarks) consistently reach a BER near 0.5, which means random guessing. This confirms that the original watermark is completely disrupted when the same watermarking method is reused for overwriting. In contrast, off-diagonal values (different methods) show very low BERs, suggesting that overwriting using a different watermarking method fails to destroy the original legitimate watermark. This is because different watermarking methods operate in distinct embedding domains and rely on method-specific decoding mechanisms.
Metric Timbre AudioSeal WavMark
ASR 99.80 100.00 100.00
ACC 100.00 100.00 100.00
@@
In the gray-box setting, the adversary knows the watermarking system’s architecture but lacks access to its training details or training data. To assess this scenario, we construct surrogate models under two settings: (i) cross-training, where the surrogate model is trained on the same dataset (VoxCeleb1) but with a different training pipeline, and (ii) cross-data, where the surrogate is trained on an entirely different dataset (LibriSpeech). We train three surrogate models with varying random seeds, denoted as Init-1, Init-2, and Init-3, respectively. The “Init-$k\rightarrowOfficial$” notation indicates that surrogate Init-$k$ attempts to overwrite the watermark embedded by official model.
@@
Method (Libri $\rightarrow$ Vox) Init-1 Init-2 Init-3
@@
The query-based attack mitigates these drawbacks by utilizing limited queries to the watermark detector. Instead of exhaustively applying all candidates, the adversary iteratively “try-and-test” candidates and stops once the detector confirms successful overwriting of the original legitimate watermark. Consequently, audio degradation is limited to a single embedding operation. As depicted in Figure 5, watermarking models exhibit varying convergence time. Certain methods reach sufficient overwriting capabilities early in the training, allowing effective overwriting with partially trained embedders. Table 4 provides a comparative analysis between zero-query and query-based attacks. The results indicate that query-based attacks, using fewer than 10 detector queries, achieve a reduction of over 50% in training iterations compared to the zero-query approach. By applying only a single effective watermarking method instead of stacking multiple methods, query-based attacks preserve the audio quality and maintain identical attack success rates. As the size of the candidate set grows, the advantages of the query-based approach become more pronounced.
Attack Type Query Training Cost SNR (dB) ASR (%)
Zero-query 0 36,000 iters 20.63 100
Query-based <<10 14,000 iters 24.19 100
@@
Acknowledgements
The work of L. Yao, C. Huang, and M. Pan was supported in part by the US National Science Foundation under grants CNS-2107057, CNS-2318664, CSR-2403249, and CNS-2431596. The work of P. Lin was supported in part by the National Science and Technology Council (NSTC) of Taiwan under grants NSTC 114- 2221-E-002-141-MY3, NSTC 112-2221-E-002-163-MY3, and NSTC 113-2314-B-A49-034-MY3. The work of T. Ohtsuki was supported in part by JST ASPIRE Grant Number JPMJAP2326, Japan.